# Emergence of Social Norms in Large Language Model-based Agent Societies

### v1

@@ Abstract
22footnotetext: Co-first, equal contributions11footnotetext: Corresponding authorsThe emergence of social norms has attracted much interest in a wide array of disciplines, ranging from social science and cognitive science to artificial intelligence. In this paper, we propose the first generative agent architecture that empowers the emergence of social norms within a population of large language model-based agents. Our architecture, named CRSEC, consists of four modules: Creation & Representation, Spreading, Evaluation, and Compliance. Our architecture addresses several important aspects of the emergent processes all in one: (i) where social norms come from, (ii) how they are formally represented, (iii) how they spread through agents’ communications and observations, (iv) how they are examined with a sanity check and synthesized in the long term, and (v) how they are incorporated into agents’ planning and actions. Our experiments deployed in the Smallville sandbox game environment demonstrate the capability of our architecture to establish social norms and reduce social conflicts within large language model-based multi-agent systems. The positive outcomes of our human evaluation, conducted with 30 evaluators, further affirm the effectiveness of our approach.
@@ 1 Introduction
Social norms are standards of behavior shared within a social group [53]. In human societies, social norms have shaped almost every aspect of our daily life, from the language we speak and the etiquette we drive to the amount we tip. Without social norms, we may feel confused about how to behave appropriately in social situations and, consequently, social conflicts may arise [30]. Over the past decades, the study of social norms has attracted much interest in a variety of disciplines, such as economics [61], cognitive science [25], computational social science [7], complex system science [14], and computer science [40]. Across these studies, a central question is: how do social norms spontaneously emerge from social interactions of humans or agents?
@@
The deployment of AI systems, particularly multi-agent systems (MASs), in real-world situations raises the need for these systems to be normative—the capability of empowering agents to understand certain standards of behavior and behave appropriately according to the standards [11, 20]. Imagine that agents within a system interact with other agents or humans to accomplish some tasks; for the system to be truly accepted and embraced by humans, such a capability will be crucial, as it can reduce conflicts within systems, enable more effective coordination among agents (potentially including humans), and even allow humans to better anticipate the system’s behaviors—a key means to improve human trust in the system [6, 4, 8, 16].
@@
This paper studies the emergence of social norms within a system of generative agents, i.e. agents that are powered by large-language models (LLMs). Recently, LLMs, such as GPT, BERT, and Gemini, have undoubtedly taken the world by storm. Since LLMs are trained on extensive corpora of human text, it is not surprising that they may inherently embed social norms [49, 23]. It could be questioned then what is the point of fostering the emergence of social norms within generative MASs? While LLMs can capture social norms, it has also been shown that LLMs do not adequately understand social norms, especially the culture-specific ones [44, 24]. This deficiency can provoke conflicts among generative agents, particularly when their base LLMs are trained on text corpora from diverse cultural backgrounds. Moreover, as generative agents increasingly become more personalized (such as functioning as personal assistants) and represent humans in social situations, it is natural to expect that these agents, reflecting the values and preferences of their human users, will encounter social conflicts similar to those experienced by humans. To tackle these challenges, approaches must go beyond merely embedding LLMs with human norms or aligning them to such norms [34, 32]; rather, they should be able to foster the emergence of social norms within generative MASs so that generative agents can establish their own standards of behavior out of their interactions and adhere to these standards to address those conflicts.
@@
How can we empower generative MASs with the capability to foster social norm emergence? We argue that the key is to instigate an emergent process—generative agents, starting with initially only a few adopting certain standards of behavior, influence others and propagate these standards, ultimately resulting in widespread acceptance and adherence of these standards across the system. Recent work has shown that generative MASs can reproduce believable social behaviors (such as spreading invitations to a party and asking each other out) [43], achieve multi-agent cooperation surpassing conventional methods [62, 15], and collaboratively solve complex tasks (such as automatic code generation) [31, 27]. While these systems have demonstrated the potential of leveraging LLMs in MAS research, the emergence of social norms remains largely unaddressed in existing studies, primarily because they typically focused on fully cooperative tasks—where agents’ values, preferences, or objectives align, thereby generally preventing social conflicts and voiding the need for social norms. Fortunately, the extensive and multidisciplinary literature on social norm emergence can offer a wealth of resources for inspiration. That said, those studies cannot provide a direct solution for two reasons. First, historically, they have not been able to harness the strength of LLMs. Second, those studies often focused on just one or two aspects of the emergent process,111Similar arguments were made in surveys [48, 26]. For example, Savarimuthu and Cranefield [47] proposed the concept of norm life cycle and categorized existing studies into usually one phase of the cycle. and a tangible implementation that integrates various aspects has been missing; for example, while some studies may focus on norm representation [21, 3], others may delve into norm compliance and enforcement [37, 56, 35], or explore norm learning [51, 9, 29, 28].
@@
In this paper, we propose, to our knowledge, the first normative architecture for generative MASs. This architecture, which can be abbreviated as CRSEC, consists of four modules: Creation & Representation, Spreading, Evaluation, and Compliance. Through the Creation & Representation module, norm entrepreneurs (agents who actively campaign norms) can generate their own personal standards of behavior (or personal norms), and these standards are formally represented and stored in their databases. Through the Spreading module, some agents influence others to adopt the standards via communication, while others can identify these standards by reflecting on their conversations and observations. With the Evaluation module, agents perform a sanity check to decide whether they accept certain standards as their own personal norms, and, from time to time, synthesize their personal norms to keep the norms compact and concise. Lastly, the Compliance module raises agents’ awareness of their personal norms, encouraging them to generate plans and take actions in line with the norms. An overview of our architecture is shown in Figure 1.
@@
To see if and how our architecture leads to the emergence of social norms within generative MASs, we ran our experiments on the Smallville sandbox game environment [43], and simulated the scenarios where initially agents have conflicts in their values and preferences. We show that social norms always emerge in multiple independent runs of our experiments, leading to 100% of agents accepting some standards of behaviors as their personal norms and complying with these norms in their plans and actions; moreover, as social norms emerge, social conflicts almost vanish. Moreover, we observe that conversations and thoughts drive the emergence of social norms, and descriptive norms are harder to establish than injunctive norms, yet norm entrepreneurs can shape their emergence. For a better understanding, we additionally present a case study to illustrate how a seasoned smoker in Smallville’s environment has been gradually persuaded to accept “no smoking indoors” as his own personal norm, and eventually even stepped forward to remind another agent upon noticing that agent’s breach of the norm. Finally, we present the results of our human evaluation, which involved 30 evaluators, to gauge the effectiveness of our architecture from a human perspective. The feedback gathered from the questionnaires reflects an overall positive evaluation. Furthermore, interviews conducted after the questionnaires shed light on aspects that humans consider important for the emergence of social norms and suggest potential directions for future work.
@@ 2 Principles and Architecture
In this paper, we aim to design an agent architecture that facilitates the emergence of social norms within a generative agent society. Specifically, this emergent process should start with a minority of agents campaigning certain standards of behavior; then over time, these standards should be able to spread within the society, where each member, a LLM-based agent, can perceive the environment (e.g. by observation), interact with other agents (e.g. by communication), generate their own plans, and take actions; ultimately, adherence to these standards should become predominant even if it may be against personal interests, leading to the emergence of the corresponding social norm in the society. Our CRSEC architecture, which plays a pivotal role in facilitating this process, comprises four modules: Creation & Representation, Spreading, Evaluation, and Compliance modules. We detail each module in the following sections.
@@ 2.1 Creation and Representation
The Creation and Representation module of our architecture addresses the questions of where social norms come from and how they can be formally represented. In human society, social norms are usually shaped by norm entrepreneurs who actively influence and persuade others to alter their behaviors in accordance with the entrepreneurs’ personal standards of behavior (or personal norms) [55]. Personal norms typically flow from one’s values [50], and would become social norms if they were to be widely adopted by other members of a social group. According to [17], there are two types of (personal or social) norms: (i) descriptive ones that reflect what most people typically do in a given situation, and (ii) injunctive ones that dictate what ought or ought not to be done in a given situation. For example, the common practice of shaking hands upon meeting someone is descriptive; in contrast, no smoking indoors is injunctive.
@@
A distinct feature of generative agents is that by using natural language that mimics how one typically describes humans, these agents can exhibit characteristics and personalities in alignment with the agent description [52]. Recall that agents’ values or preferences typically vary, as we analyze in the introduction. To ensure that the created personal norms are consistent with norm entrepreneurs’ agent descriptions, we instruct LLMs through prompts to create these norms based on norm entrepreneurs’ agent descriptions. Let denote an agent description, and denote a set of created personal norms. We represent this LLM-based operation by . This operation not only generates personal norms in natural language, but also classifies a newly formed personal norm (i.e., deciding the value of ), and also assesses the utility of that norm on a scale of to . Once created, personal norms are stored in each norm entrepreneur’s personal norm database.
@@ 2.2.1 Communication between Agents
Generative agents are well known for their capability to generate human-like conversations [18]. It is thus natural to consider spreading norms by leveraging such a capability. To achieve this, we consider two perspectives: a sender’s perspective and a receiver’s perspective.
@@
In human societies, the desire to resolve social conflicts has driven the emergence of numerous social norms [12]. Inspired by this, we instruct each generative agent (a sender) to detect if there are any observations of other agents’ behaviors that conflict with its personal norms. Let be the text description of the sender’s observations of the environment, and be the sender’s set of qualified personal norms in its database. We represent this LLM-based operation by . Note that initially, since only norm entrepreneurs have their own personal norms, ordinary agents will detect no conflicts. However, as time evolves, ordinary agents may also develop their personal norms, and thus conflicts may be detected. Once a conflict is detected (i.e., ), then the sender will decide whether to proactively step in and start a conversation in order to influence others and propagate its personal norms. Intuitively, if the sender is a norm entrepreneur, then it will start a conversation without doubt, as it is interested in influencing others. However, if the sender is an introverted, ordinary agent, it may not start a conversation. Thus, for better autonomy, we let the sender decide based on its agent description , and represent this LLM-based operation by .
@@ 2.2.2 Observation from Others’ Behavior
Observation has long been recognized as a key mechanism for humans and agents to learn norms [41, 54, 10, 42]. Recent work has shown that generative agents can generate thoughts from the text description of their observations [43, 33]. Let denote the text description of observations, and denote the generated thoughts. This LLM-based operation can be denoted by , and it can be achieved by modules that generate thoughts in existing studies. Leveraging on this, we prompt generative agents to identify normative information from the generated thoughts. We represent this LLM-based operation by , where represents normative information. Note that this operation is similar to , as both these operations can be viewed as a kind of text summarizing tasks. Once generated, the normative information is also stored in the personal norm database with the deactivated and invalid state ().
@@ 2.3.2 Long-term Synthesis
Over time, as agents accumulate more qualified personal norms, they accept a broader range of standards of behavior, which could potentially limit their liberty. Morales et al. [38, 39] suggested that for better agent liberty, it would be beneficial to synthesize norms into a compact and concise set of possibly more abstract ones. Inspired by this, we prompt each generative agent to start a synthesis within its personal norm database if the sum of the utility of its qualified personal norms exceeds a certain threshold.222One can, of course, change the condition for starting the synthesis. In general, the principle is that this synthesis takes effect only after agents have already accumulated plentiful personal norms.
@@ 2.4 Compliance
The Compliance module of our architecture raises agents’ awareness of personal norms in their behaviors. Note that with such an awareness, agents can choose to comply with the norms or not, thereby granting them with greater autonomy [19, 20]. We design this module focusing on two sub-components: (i) compliance in planning, and (ii) compliance in actions.
@@ 2.4.1 Compliance in Planning
Plans describe a sequence of actions for agents. Recent work has shown that generative agents are good at planning towards some goals; this ensures that agents’ behaviors are consistent over time [57, 33]. Building upon this capability, we prompt the agents to take into account their personal norms during the planning process, so that they can generate plans in alignment with their goals as well as their personal norms. Let denote a plan (e.g. 10:30 am to 11:00 am: Have a light breakfast), and denote a list of plans (e.g. for every hour in a day). The planning process of our architecture can be represented by , where the inputs are the current goals and the set of qualified personal norms.
@@ 3.1 Experimental Settings
Our experiments were conducted in Park et al. [43]’s Smallville sandbox game environment, which is arguably the most well-known environment for generative MASs. This environment offers a variety of scenarios where LLM-based agents can exhibit human-like behaviors, including observation, interaction with others, planning, and action execution. In our setup, there were 10 generative agents, including 3 norm entrepreneurs and 7 ordinary agents. To simulate scenarios where individuals can have conflicts in values or preferences, we considered that ordinary agents’ agent descriptions exhibited diverse inclinations: some favored smoking in public, speaking loudly, or supporting a tipping culture, whereas others held opposite preferences. For norm entrepreneurs’ agent descriptions, we considered all of them to favor “no smoking indoors” and “be quiet in public”. However, since whether to tip varies across cultures, we considered two of them to support tipping while one did not. In addition to these preferences, each agent’s agent description also included its name, personality, occupation, short-term goal, and social relationships with other agents, etc. Details of agent descriptions and experimental parameters, such as the number of initial personal norms for norm entrepreneurs and the threshold for starting a synthesis in the Evaluation module, are provided in Appendix A.
Our implementation utilized GPT-3.5 and GPT-4. Using the same experimental setup, we repeated our experiments for 5 runs. To be time-efficient and cost-efficient, we focused on the scenario “Hobbs Café” (as visualized at the bottom left corner of Figure 1) and let the experiments continue for 2 days in the Smallville environment. That said, each run cost more than $500 dollars and about 7 days to complete. The GitHub repository for our project can be accessed via the following link: https://github.com/sxswz213/CRSEC.

### v2

# Emergence of Social Norms in Generative Agent Societies: Principles and Architecture

@@ Abstract
22footnotetext: Co-first, equal contributions11footnotetext: Corresponding authorsSocial norms play a crucial role in guiding agents towards understanding and adhering to standards of behavior, thus reducing social conflicts within multi-agent systems (MASs). However, current LLM-based (or generative) MASs lack the capability to be normative. In this paper, we propose a novel architecture, named CRSEC, to empower the emergence of social norms within generative MASs. Our architecture consists of four modules: Creation & Representation, Spreading, Evaluation, and Compliance. This addresses several important aspects of the emergent processes all in one: (i) where social norms come from, (ii) how they are formally represented, (iii) how they spread through agents’ communications and observations, (iv) how they are examined with a sanity check and synthesized in the long term, and (v) how they are incorporated into agents’ planning and actions. Our experiments deployed in the Smallville sandbox game environment demonstrate the capability of our architecture to establish social norms and reduce social conflicts within generative MASs. The positive outcomes of our human evaluation, conducted with 30 evaluators, further affirm the effectiveness of our approach. Our project can be accessed via the following link: https://github.com/sxswz213/CRSEC.
@@ 1 Introduction
In human societies, social norms, which are standards of behavior shared within a social group Sherif (1936), have shaped almost every aspect of our daily life, from the language we speak and the etiquette we drive to the amount we tip. Without social norms, people may feel confused about how to behave appropriately in social situations and consequently social conflicts may arise Lewis (1969). Over the past decades, the study of social norms has attracted much interest in a variety of disciplines, such as economics Young (2015), cognitive science Hawkins et al. (2019), complex system science Centola et al. (2018), and computer science Morris-Martin et al. (2019). Across these studies, a central question is: how do social norms spontaneously emerge from social interactions of humans or agents?
@@
This paper studies the emergence of social norms within a generative multi-agent system (MAS), i.e. a system of agents that are powered by large-language models (LLMs). The deployment of MASs in real-world situations raises the need for these systems to be normative—the capability of empowering agents to understand certain standards of behavior and behave appropriately according to the standards Boella et al. (2008); Criado et al. (2011). Imagine that agents within a system interact with other agents or humans to accomplish some tasks; for the system to be truly accepted and embraced by humans, such a capability will be crucial, as it can reduce conflicts within systems, enable more effective coordination among agents (potentially including humans), and allow humans to anticipate the system’s behaviors—a key means to improve human trust in the system Awad et al. (2018); Ajmeri et al. (2020); Chugunova and Sele (2022).
@@
Since LLMs are trained on extensive corpora of human text, it is not surprising that they may inherently embed social norms Schramowski et al. (2022); Guo et al. (2023). One might thus challenge the importance of fostering the emergence of social norms within generative MASs. While LLMs can capture social norms, it has also been shown that LLMs do not adequately understand social norms, especially the culture-specific ones Ramezani and Xu (2023); Hämmerl et al. (2022). This deficiency can provoke conflicts among generative agents, particularly when their base LLMs are trained on text corpora from diverse cultural backgrounds. Moreover, as generative agents increasingly become more personalized (such as functioning as personal assistants) and represent humans in social situations, it is natural to expect that these agents, reflecting the values and preferences of their human users, will encounter social conflicts similar to those experienced by humans. To tackle these challenges, approaches must go beyond merely embedding LLMs with human norms or aligning them to such norms Liu et al. (2024); Li et al. (2024); rather, they should also be able to foster the emergence of social norms within generative MASs so that generative agents can establish their own standards of behavior out of their interactions and adhere to these standards to address those conflicts.
@@
How can we empower generative MASs with the capability to foster social norm emergence? We argue that the key is to instigate an emergent process—generative agents, starting with initially only a few adopting certain standards of behavior, influence others and propagate these standards, ultimately resulting in widespread acceptance and adherence of these standards across the system. Recent work has shown that generative MASs can reproduce believable social behaviors (such as spreading invitations to a party) Park et al. (2023), achieve multi-agent cooperation surpassing conventional methods Zhang et al. (2024), and collaboratively solve complex tasks (such as automatic code generation) Hong et al. (2023). While these systems have demonstrated the potential of leveraging LLMs in MAS research, the emergence of social norms remains largely unaddressed in existing studies, primarily because they typically focused on fully cooperative tasks—where agents’ values, preferences, or objectives align, thereby generally preventing social conflicts and voiding the need for social norms.
Fortunately, the extensive and multidisciplinary literature on social norm emergence can offer a wealth of resources for inspiration. For example, some studies may focus on norm representation Dignum (1999); Ågotnes et al. (2009), some may delve into norm compliance and enforcement Modgil et al. (2009); Villatoro et al. (2011); Mahmoud et al. (2015), and others may explore norm learning Sen and Airiau (2007); Beheshti et al. (2015); Hu and Leung (2017); Hu et al. (2019). That said, these studies cannot provide a direct solution for two key reasons. First, historically, they have not been able to harness the strength of LLMs. Second, they typically focused on isolated aspects of the emergent process, consequently leaving the tangible implementation that integrates various aspects as an open problem Savarimuthu and Cranefield (2011); Haynes et al. (2017).
@@
In this paper, we propose, to our knowledge, the first normative architecture for generative MASs. Our architecture, abbreviated as CRSEC, consists of four modules: Creation & Representation, Spreading, Evaluation, and Compliance. This architecture not only fosters the emergence of social norms within generative MASs, but also addresses the open problem of actualizing various aspects of the emergent process into an operational framework. Specifically, through the Creation & Representation module, norm entrepreneurs (agents who actively campaign norms) can generate their own personal standards of behavior (or personal norms), and these standards are formally represented and stored in their databases. Through the Spreading module, some agents influence others to adopt the standards via communication and actual behaviors, while others can identify these standards by reflecting on their conversations and observations. With the Evaluation module, agents perform a sanity check to decide whether they accept certain standards as their own personal norms, and, from time to time, synthesize their personal norms to keep the norms compact and concise. Lastly, the Compliance module raises agents’ awareness of their personal norms, encouraging them to generate plans and take actions in line with the norms. An overview of our architecture is shown in Figure 1.
@@
To verify if and how our architecture leads to the emergence of social norms within generative MASs, we ran our experiments on the Smallville sandbox game environment Park et al. (2023), and simulated the scenarios where initially agents have conflicts in their values and preferences. We show that social norms always emerge in multiple independent runs of our experiments, leading to 100% of agents accepting some standards of behaviors as their personal norms and complying with these norms in their plans and actions; moreover, as social norms emerge, social conflicts almost vanish. Moreover, we observe that conversations and thoughts drive the emergence of social norms, and descriptive norms are harder to establish than injunctive norms, yet norm entrepreneurs can shape their emergence. For a better understanding, we additionally present a case study to illustrate how a seasoned smoker in Smallville’s environment has been gradually persuaded to accept “no smoking indoors” as his own personal norm, and eventually even stepped forward to remind another agent upon noticing that agent’s breach of the norm. Finally, we present the results of our human evaluation, which involved 30 evaluators, to gauge the effectiveness of our architecture from a human perspective. The feedback gathered from the questionnaires reflects an overall positive evaluation. Furthermore, interviews conducted after the questionnaires shed light on aspects that humans consider important for the emergence of social norms and suggest potential directions for future work.
@@ 2.1 Creation and Representation
The Creation and Representation module of our architecture addresses the questions of where social norms come from and how they can be formally represented. In human society, social norms are usually shaped by norm entrepreneurs, who actively influence and persuade others to alter their behaviors in accordance with the entrepreneurs’ personal standards of behavior (or personal norms) Sunstein (1996). Personal norms typically flow from one’s values Schwartz (1973), and would become social norms if they were to be widely adopted by other members of a social group. According to Cialdini et al. (1991), there are two types of (personal or social) norms: (i) descriptive ones that reflect what most people typically do in a given situation, and (ii) injunctive ones that dictate what ought or ought not to be done in a given situation. For example, the common practice of shaking hands upon meeting someone is descriptive; in contrast, no smoking indoors is injunctive.
@@
A distinct feature of generative agents is that by using natural language that mimics how one typically describes humans, these agents can exhibit characteristics and personalities in alignment with the agent description Shanahan et al. (2023). Recall that agents’ values or preferences typically vary, as we analyze in the introduction. To ensure that the created personal norms are consistent with norm entrepreneurs’ agent descriptions, we instruct LLMs through prompts to create these norms based on norm entrepreneurs’ agent descriptions. Let denote an agent description, and denote a set of created personal norms. We represent this LLM-based operation by . This operation not only generates personal norms in natural language, but also classifies a newly formed personal norm (i.e., deciding the value of ), and also assesses the utility of that norm on a scale of to . Once created, personal norms are stored in each norm entrepreneur’s personal norm database.
@@ 2.2.1 Communication between Agents
Generative agents are well known for their capability to generate human-like conversations Clark et al. (2021). It is thus natural to consider spreading norms by leveraging such a capability. To achieve this, we consider two perspectives: a sender’s perspective and a receiver’s perspective.
@@
In human societies, the desire to resolve social conflicts has driven the emergence of numerous social norms Nyla R. Branscombe (2022). Inspired by this, we instruct each generative agent (a sender) to detect if there are any observations of other agents’ behaviors that conflict with its personal norms. Let be the text description of the sender’s observations of the environment, and be the sender’s set of qualified personal norms in its database. We represent this LLM-based operation by . Note that initially, since only norm entrepreneurs have their own personal norms, ordinary agents will detect no conflicts. However, as time evolves, ordinary agents may also develop their personal norms, and thus conflicts may be detected. Once a conflict is detected (i.e., ), then the sender will decide whether to proactively step in and start a conversation in order to influence others and propagate its personal norms. Intuitively, if the sender is a norm entrepreneur, then it will start a conversation without doubt, as it is interested in influencing others. However, if the sender is an introverted, ordinary agent, it may not start a conversation. Thus, for better autonomy, we let the sender decide based on its agent description , and represent this LLM-based operation by .
@@ 2.2.2 Observation from Others’ Behavior
Observation has long been recognized as a key mechanism for humans and agents to learn norms Nakamaru and Levin (2004); Shettleworth (2009); Beheshti and Sukthankar (2014); Paiva et al. (2018). Recent work has shown that generative agents can generate thoughts from the text description of their observations Park et al. (2023); Lin et al. (2023). Let denote the text description of observations, and denote the generated thoughts. This LLM-based operation can be denoted by , and it can be achieved by modules that generate thoughts in existing studies. Leveraging on this, we prompt generative agents to identify normative information from the generated thoughts. We represent this LLM-based operation by ,where represents normative information. Note that this operation is similar to , as both these operations can be viewed as a kind of text summarizing tasks. Once generated, the normative information is also stored in the personal norm database with the deactivated and invalid state ().
@@ 2.3.2 Long-term Synthesis
Over time, as agents accumulate more qualified personal norms, they accept a broader range of standards of behavior, which could potentially limit their liberty. Morales et al. Morales et al. (2013, 2015) suggested that for better agent liberty, it would be beneficial to synthesize norms into a compact and concise set of possibly more abstract ones. Inspired by this, we prompt each generative agent to start a synthesis within its personal norm database if the sum of the utility of its qualified personal norms exceeds a certain threshold.
@@ 2.4 Compliance
The Compliance module of our architecture raises agents’ awareness of personal norms in their behaviors. Note that with such an awareness, agents can choose to comply with the norms or not, thereby granting them with greater autonomy Conte et al. (1998); Criado et al. (2011). We design this module focusing on two subcomponents: (i) compliance in planning, and (ii) compliance in actions.
@@ 2.4.1 Compliance in Planning
Plans describe a sequence of actions for agents. Recent work has shown that generative agents are good at planning towards some goals; this ensures that agents’ behaviors are consistent over time Wang et al. (2023); Lin et al. (2023). Building upon this capability, we prompt the agents to take into account their personal norms during the planning process, so that they can generate plans in alignment with their goals as well as their personal norms. Let denote a plan (e.g. 10:30 am to 11:00 am: Have a light breakfast), and denote a list of plans (e.g. for every hour in a day). The planning process of our architecture can be represented by , where the inputs are the current goals and the set of qualified personal norms.
@@ 3.1 Experimental Settings
Our experiments were conducted in Park et al. Park et al. (2023)’s Smallville sandbox game environment, which is arguably the most well-known environment for generative MASs. This environment offers a variety of scenarios where LLM-based agents can exhibit human-like behaviors, including observation, interaction with others, planning, and action execution. In our setup, there were 10 generative agents, including 3 norm entrepreneurs and 7 ordinary agents. To simulate scenarios where individuals can have conflicts in values or preferences, we considered that ordinary agents’ agent descriptions exhibited diverse inclinations: some favored smoking in public, speaking loudly, or supporting a tipping culture, whereas others held opposite preferences. For norm entrepreneurs’ agent descriptions, we considered all of them to favor “no smoking indoors” and “be quiet in public”. However, since whether to tip varies across cultures, we considered two of them to support tipping while one did not. In addition to these preferences, each agent’s agent description also included its name, personality, occupation, short-term goal, and social relationships with other agents, etc. Details of agent descriptions and experimental parameters, such as the number of initial personal norms and the threshold for starting a synthesis in the Evaluation module, are provided in Appendix A.
Our implementation utilized GPT-3.5 and GPT-4. Using the same experimental setup, we repeated our experiments for 5 runs. To be time-efficient and cost-efficient, we focused on the scenario “Hobbs Café” (as visualized at the bottom left corner of Figure 1) and let the experiments continue for 2 days in the Smallville environment. Each run costs more than $500 dollars and about 7 days to complete. The GitHub repository for our project can be accessed via the following link: https://github.com/sxswz213/CRSEC.